What is the change in energy (in joules) if a reaction loses 399 J of heat and decreases in volume from 0.650 L to 0.225 L at a constant pressure of 4.65 atm? Please include the correct sign with your numerical result.
Answer #1


E = q + w

where E = change in energy

q = heat

w = work

work done by the system (w) = -p *v

                                                   = - 4.65 atm * (0.650-0.225)L

                                                  = -1.97625 L.atm

but 1 L atm---------------------------> 101.3 joules

1.97625 L atm ----------------------> ?

=> 1.97625*101.3 =   200.2 joules

w=> 200.2 joules

q = 399 joules

if work done by the system then

E = q + w

       = 399 + 200.2

       = 599.2 joules